OpenStudy (anonymous):

lve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

1. Open brackets. 2. Take all terms containing x to one side. 3. Solve. 3x + 3 = -2x + 2 - 4 3x + 2x = 2 - 4 - 3 5x = 2 - 7 5x = -5 x = -5/5 = -1

OpenStudy (anonymous):

ahh thank you so much!(: